OpenStudy (anonymous):

Hi guys I have a math assignment, and I am having trouble understanding it. I know how to solve it, but it's asking for a function, here it is "Create your own function to teach the Martians about functions. Your function must contain at least two different operations." This is the problem 2x+4=-3x+14 I know the answer is 2, but I don't understand the question right. Please help.

OpenStudy (phi):

a "function" is a rule. You are given a number, and the "rule" tells you what do to that number to get a "new number" Here is a *simple* example: f(x)= x the rule says, "give me a number, and the "new" number is just what we started with" (not much of a rule, but that's ok) in your case they want you to do at least 2 things to it. i.e. multiply , divide, add , subtract , exponent , square root, etc. I would stick to multiply and add

OpenStudy (phi):

your ***2x+4=-3x+14*** is an equation, which is a different idea if you said f(x) (this is the "name" of the function" uses the rule 2x+4 i.e. f(x)= 2x + 4 that would be a good answer

OpenStudy (anonymous):

OK thank you, the next part is this "Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function.". I took intermediate algebra(at a college) so I barley remember this math...... How would you plug in 2x+4? Or like where? Grr

OpenStudy (phi):

a function is a relation between input and output, with the requirement that each input is related to *exactly one* output. in other words, if you "put in" the same value, you always get back the same output value. Here is *not* a function \[ y = \pm \sqrt{x} \] if you put in 1 for x, y can be either +1 or -1. A function would allow only 1 fixed value. your equation is the equation of a line, so it is a function. if you pick an x, there will be only one y value that you get for that x value.

OpenStudy (anonymous):

Thank you a lot. I reread an older lesson, and I think I can complete this. Its somehow so confusing. grr but anyways thanks

OpenStudy (phi):

For background see these videos https://www.khanacademy.org/math/algebra2/functions_and_graphs and in particular, here is an example: https://www.khanacademy.org/math/algebra2/functions_and_graphs/recognizing-functions-2/v/recognizing-functions-example-1 there are also other examples.

OpenStudy (anonymous):

Sorry, I forgot to mark you as best response. But I figured something out :)